Examples: (10 points each) . A small country is unable to affect world price. It imports peanuts at the price of USD 20 per bag Demand curve D 450 -8P Supply curve: S-20+4P a) Determine the free trade equilibrium (the amount of import under free trade) b) Calculate the effect of an import quota that limits imports to 100 bags
Answer #1

Ans. a) In order to determine the free trade equilibrium , we have to substitute the P = 20 into demand and supply schedule.

D = 450 - 8(20) = 290

S = 20 + 4(20) = 100

Total Imports = 290 - 100 = 190

b) Effect of Import Quota on Price would be as follows :

D = S + Quota

450 - 8P = 20 + 4P + 100

450 - 120 = 12P

330 = 12P

P = 27.5

Best of Luck !! !!
Answer #2
400-10(p) 400-10(10) D=300 50+5(p) 50+5(5) 100
answered by: gak_i27